# OPERA: A Reinforcement Learning–Enhanced Orchestrated Planner-Executor Architecture for Reasoning-Oriented Multi-Hop Retrieval

### v1

@@ Acknowledgments
This work is supported by the National Key R&D Program of China (No. 2023YFC3303800), National Natural Science Foundation of China (No. 62406161), China Postdoctoral Science Foundation (No. 2023M741950), and the Postdoctoral Fellowship Program of CPSF (No. GZB20230347).
@@ Appendix A Appendix
Table 6: MAPGRPO training hyperparameters for each agent.
@@
aQwen2.5 series (Yang et al. 2024). bLLama3.1 series (Grattafiori et al. 2024). cQwen3 (Yang et al. 2025).
@@
Table 11: Error distribution analysis across agent behaviors.
@@
Table 12: Success patterns in correctly answered questions.
@@
Planning Inefficiencies. Many questions allow multiple valid decomposition paths, but OPERA lacks explicit optimization for path efficiency and may select suboptimal reasoning chains. Several research directions could address these limitations. Developing adaptive decomposition methods with path-efficiency optimization would improve planning effectiveness. Creating specialized datasets for long-chain reasoning would enhance performance on complex multi-hop questions. Hybrid architectures that maintain modularity while reducing computational overhead could make the system more practical for real-time use. Advanced retrieval techniques like Beam Retrieval (Zhang et al. 2024) could explore multiple retrieval paths simultaneously, improving robustness for ambiguous queries.

### v2

@@ Acknowledgments
This research is supported by the National Key R&D Program of China through grant 2023YFC3303800 and Procurement Project through grant E5V01511D3, the National Natural Science Foundation of China (No. 62406161), the China Postdoctoral Science Foundation (No. 2023M741950) and the Postdoctoral Fellowship Program of CPSF (No. GZB20230347).
@@ Appendix A Appendix
Table 6: MAPGRPO training hyperparameters for each agent.
@@
aQwen2.5 series (yang2024qwen2.5). bLLama3.1 series (grattafiori2024llama3). cQwen3 (yang2025qwen3).
@@
Table 11: Error distribution analysis across agent behaviors.
@@
Table 12: Success patterns in correctly answered questions.
@@
Planning Inefficiencies. Many questions allow multiple valid decomposition paths, but OPERA lacks explicit optimization for path efficiency and may select suboptimal reasoning chains. Several research directions could address these limitations. Developing adaptive decomposition methods with path-efficiency optimization would improve planning effectiveness. Creating specialized datasets for long-chain reasoning would enhance performance on complex multi-hop questions. Hybrid architectures that maintain modularity while reducing computational overhead could make the system more practical for real-time use. Advanced retrieval techniques like Beam Retrieval (zhang2024endtoend) could explore multiple retrieval paths simultaneously, improving robustness for ambiguous queries.